# Exploring the Landscape of Non-Equilibrium Memories with Neural Cellular Automata

### v1

@@ Introduction:
The design of many-body systems that retain memory of their initial conditions in the presence of perturbations, i.e. systems which function as robust many-body memories, is a fundamental problem at the intersection of non-equilibrium statistical mechanics [1], fault-tolerant computation [2], and biological information processing [3]. Memory in locally-interacting systems must be achieved through collective, distributed error correction: errors must be autonomously detected and corrected by the noisy dynamics itself, and the general mechanisms by which this may be accomplished are only partially understood [4, 2, 5, 6, 7, 8, 9, 10, 11, 12, 13].
@@
The proof makes use of a theorem of Toom [6] regarding monotone CA (viz. those for which $\mathcal{A}$ defines a monotonic Boolean function  [^2] A function $f(x)\,:\,\{-1,1\}^{n}\rightarrow\{-1,1\}$ is monotonic if increasing the number of 1s in the input cannot change an output of 1 to an output of 1-1., such as a majority vote), which states that any monotone eroder is a robust memory at small enough noise strengths. We prove erosion by proving that minority domains are always eroded ballistically along certain directions if $\mathcal{V}_{\mathbf{r}}$ has no $C_{\pi}$ symmetry. Conversely, with $C_{\pi}$ symmetry, we show that there are always certain shapes of minority domains that, regardless of their size, are never corrected by $\mathcal{A}$; under biased noise, these domains grow and rapidly destroy the encoded information. The details are provided in App. III.
@@
As stated, theorem 1 only applies to dynamics with synchronous updates. The extension to asynchronous updates is nontrivial, as there exist systems which retain memory only when updated synchronously [18]. To establish results for asynchronous dynamics, we therefore turn to numerics. We use Monte Carlo simulations to estimate the logical failure rate, defined as the probability for a magnetization flip to occur by time $O(L)$ when the system is initialized in its least stable logical state:
$p_{\sf log}=\max_{s}\mathsf{P}[{\sf maj}(\mathcal{A}^{cL}(s))\neq{\sf maj}(s)],$ (1)
where the maximum is over states $s$ with long-lived magnetizations (e.g. the states with all $1$s or all 1-1s), and $c$ is a constant, fixed at $c=5$ in our numerics. For any memory we must have $\lim_{L\rightarrow\infty}p_{\sf log}=0$, and for the CA under consideration we expect, for small enough $p$, the scaling
@@
for constants $p_{c},\lambda$  [^3] This is due to the fact that for all of the memories considered in the main text, flipping $\Theta(2L)$ spins along two non-contractible loops that span each cycle of the torus is sufficient to cause a magnetization flip.. $p_{\sf log}$} for $\mathsf{A}$ dynamics with asynchronous updates and $\eta=0$ is plotted in Fig. 1 d, where we observe a threshold crossing at $p_{c}\approx 26\%$—significantly higher than the $p_{c,{\sf Toom}}\sim 14\%$ of Toom’s rule [20, 10].
@@
Do there exist robust memories beyond asymmetric majority voters? Addressing this question via brute-force search is impossible: the number of CA rules taking $3\times 3$ blocks of spins as input is $2^{2^{9}}$—vastly greater than the number of atoms in the observable universe—and random rules are empirically observed to have a negligibly small chance of being robust memories. To more efficiently explore the space of possible dynamics, we use a Machine Learning (ML)-guided approach inspired by the Neural Cellular Automaton (NCA) framework [8]. In this setup, $\mathcal{A}$ is initially parametrized as a random continuous-state CA rule (operating on $\mathbb{R}$-valued spins), implemented by a small neural network. The neural network is optimized using gradient descent to minimize a loss function that is designed to measure $\mathcal{A}$’s ability to erase large domains of errors in the presence of noise. Including large error regions in the training data, instead of relying on their spontaneous occurrence from noisy dynamics, enables faster and more targeted learning. After a fixed number of optimization steps, we discretize $\mathcal{A}$ to $\mathbb{Z}_{2}$-valued spins. Further details on the NCA architecture and training are provided in App. II.
In a set of 1000 training runs with different random seeds, 37 cellular automata converged to robust memories. None of these 37 memory rules were symmetric under $s_{\mathbf{r}}\mapsto-s_{\mathbf{r}}$ spin-flip, and all but 6 were monotonic. Symmetric rules occurred only outside the memory class, where they were common: out of the 387 symmetric rules observed, all were monotonic, and none were memories. Monotonicity was similarly widespread beyond the memory class, with 994 monotonic rules in total. In most cases where the learned rule was not a memory, the dynamics was observed to perform something similar to a $C_{\pi}$-symmetric majority vote. Finally, while training was performed only with synchronous updates, all of the learned memory rules were observed to remain robust after making the updates asynchronous.
@@
Table 1: Approximate thresholds (in percent) of the memories studied in the main text, for both asynchronous (top) and synchronous (bottom) updates. The three values $x/y/z$ indicate the maximum noise probability thresholds at biases of $\eta=1,0,$ and 1-1, respectively. “noise-stab” indicates that the memory is present only at nonzero noise; see the main text.
@@
Interestingly, the CA rules $\mathsf{B},\mathsf{C},\mathsf{D}$} found in training are observed to be unstable in MF: for these rules, $F_{\rm eff}$ always has a unique minimum, regardless of $p$. These are the first examples known to the authors of systems where MF theory entirely misses an ordered phase, and incorrectly predicts a disordered phase for all $p$. The order in these systems is thus fluctuation-stabilized  [^4] This phenomenon might be called “order by disorder”, but is quite distinct from what is usually meant by this phrase, which conventionally refers to situations where (fine-tuned) systems with many degenerate ground states have their degeneracy lifted by fluctuations. Conventional “order by disorder” decreases the number of degenerate states, and thus in some sense actually reduces the amount of order. By contrast, the dynamics considered in this work uses fluctuations to increase the number of steady states, thereby increasing the amount of order. , and future work [18] will show how fluctuations beyond MF stabilize the ordered phase.
@@
To further explore this phenomenon, consider updates where a fraction $\alpha\in[0,1]$ of sites are simultaneously updated at each time step. At small enough $p$, each of the studied memories undergoes a percolation transition at an intermediate update fraction $\alpha_{c}$ between a robust memory phase (at small $\alpha$), and a trivial phase dominated by $1$s (at large $\alpha$); this phenomenon is somewhat similar to the synchronicity transitions that occur in other (non-robust) CA, such as Conway’s game of life [26, 27]. As shown in Fig. 2, the spatial spin patterns produced across this transition can be quite rich.
Figure 3: Top: Probability that a rule remains a memory $P_{\sf mem}$ as a function of the number of “mutations” made to its rule table; each data point is based on sampling 5000 randomly mutated rules. Bottom: Hamming distances between the rule tables of the learned NCA and majority-vote memories.
@@
The phenomena discovered in this work indicate that the physics of non-equilibrium memories is quite diverse, and far from fully understood. Immediate future reserach directions include generalizing the NCA framework to $\mathbb{Z}_{N}$ and $U(1)$ memories, investigating the phase transitions that occur when the memory is lost, and theoretically characterizing the distinct universality classes of local error-correcting dynamics.
We close with some rather speculative remarks. The systems studied in this work are capable of “homeostasis”: the robust protection of (here, a single bit of) information against degredation by noise. It is natural to wonder about mechanisms by which richer behavior, like nontrivial information processing and environmental adaptation, can robustly emerge in locally-interacting systems. It is well known that simple CA rules can display remarkably life-like emergent behavior [28, 29, 8], but in none of these cases has the problem of noise resilience been carefully addressed, and an outstanding problem for theory is to identify the complexity threshold that must be crossed in order for CA to exhibit complex and robust emergent behavior. The constructions of Gacs [2, 30]—whose design calls to mind the genetic code of an elaborate artificial organism—provide extremely complicated starting points, but there is clearly much more to be done. We expect the NCA approach developed in the present work to be useful in this regard.
@@ Acknowledgments:
We thank Shankar Balasubramanian, Sarang Gopalakrishnan, Vedika Khemani, Yaodong Li, Tibor Rakovsky, and Mike Zaletel for discussions and feedback. E.L. thanks Nathaniel Selub for discussions and collaboration on a related project. E.L. was supported by a Miller Research Fellowship.

### v2

@@ Introduction:
The design of many-body systems that retain memory of their initial conditions in the presence of perturbations, i.e., systems which function as robust many-body memories, is a fundamental problem at the intersection of non-equilibrium statistical mechanics [1], fault-tolerant computation [2], and biological information processing [3]. Memory in locally-interacting systems must be achieved through collective, distributed error correction: errors must be autonomously detected and corrected by the noisy dynamics itself, and the general mechanisms by which this may be accomplished are only partially understood [4, 2, 5, 6, 7, 8, 9, 10, 11, 12, 13].
@@
The proof makes use of a theorem of Toom [6] regarding monotone cellular automata (viz. those for which $\mathcal{A}$ defines a monotonic Boolean function  [^2] A function $f(x)\,:\,\{-1,1\}^{n}\rightarrow\{-1,1\}$ is monotonic if increasing the number of 1s in the input cannot change an output of 1 to an output of 1-1., such as a majority vote), which states that any monotone eroder is a robust memory at small enough noise strengths. We prove erosion by proving that minority domains are always eroded ballistically along certain directions if $\mathcal{V}_{\mathbf{r}}$ has no $C_{\pi}$ symmetry. Conversely, with $C_{\pi}$ symmetry, we show that there are always certain shapes of minority domains that, regardless of their size, are never corrected by $\mathcal{A}$; under biased noise, these domains grow and rapidly destroy the encoded information. The details are provided in App. III.
@@
As stated, Theorem 1 only applies to dynamics with synchronous updates. The extension to asynchronous updates is nontrivial, as there exist systems which retain memory only when updated synchronously [18]. To establish results for asynchronous dynamics, we therefore turn to numerics. We use Monte Carlo simulations to estimate the logical failure rate, defined as the probability for a magnetization flip to occur by time $O(L)$ when the system is initialized in its least stable memory state:
$p_{\sf log}=\max_{s_{\pm}}\mathsf{P}[{\sf maj}(\mathcal{A}^{cL}(s))\neq{\sf maj}(s)],$ (1)
where $s_{\pm}$ are respectively the all-$1$s and all-(1)(-1)s states, and $c$ is a constant, fixed at $c=5$ in our numerics. For any memory we must have $\lim_{L\rightarrow\infty}p_{\sf log}=0$, and for the CA under consideration we expect, for small enough $p$, the scaling
@@
for constants $p_{c},\lambda$  [^3] This is due to the fact that for all of the memories considered in the main text, flipping $\Theta(2L)$ spins along two non-contractible loops that span each cycle of the torus is sufficient to cause a magnetization flip.. The logical failure rate $p_{\sf log}$} for $\mathsf{A}$ dynamics with asynchronous updates and bias $\eta=0$ is plotted in Fig. 1 d, where we observe a threshold crossing at $p_{c}\approx 26\%$—significantly higher than the $p_{c,{\sf Toom}}\sim 14\%$ of Toom’s rule [20, 10].
@@
Do there exist robust memories beyond asymmetric majority voters? Addressing this question via brute-force search is impossible: the number of CA rules taking $3\times 3$ blocks of spins as input is $2^{2^{9}}$—vastly greater than the number of atoms in the observable universe—and random rules are empirically observed to have a negligibly small chance of being robust memories. To more efficiently explore the space of possible dynamics, we use a machine-learning-guided approach inspired by the neural cellular automaton (NCA) framework [8]. Using this setup, we find robust memories beyond asymmetric majority voters. In this setting, $\mathcal{A}$ is initially parameterized as a random continuous-state CA rule (operating on $\mathbb{R}$-valued spins), implemented by a small neural network. The neural network is optimized using gradient descent to minimize a loss function that is designed to measure $\mathcal{A}$’s ability to erase large domains of errors in the presence of noise. Including large error regions in the training data, instead of relying on their spontaneous occurrence from noisy dynamics, enables faster and more targeted learning. After a fixed number of optimization steps, we discretize $\mathcal{A}$ to $\mathbb{Z}_{2}$-valued spins. Further details on the NCA architecture and training are provided in App. II.
[…]
@@
Table 1: Approximate thresholds (in percent) of the memories studied in the main text, for both asynchronous (top) and synchronous (bottom) updates. The three values $x/y/z$ indicate the maximum noise probability thresholds at biases of $\eta=1,0,$ and 1-1, respectively. “noise stab.” indicates that the memory is present only at nonzero noise; see the main text.
@@
Interestingly, the CA rules $\mathsf{B},\mathsf{C},\mathsf{D}$} found in training are observed to be unstable in MF: for these rules, $F_{\rm eff}$ always has a unique minimum, regardless of $p$. These are the first examples known to the authors of systems where MF theory entirely misses an ordered phase, and incorrectly predicts a disordered phase for all $p$. The order in these systems is thus fluctuation-stabilized  [^5] This phenomenon might be called “order by disorder,” but is quite distinct from what is usually meant by this phrase, which conventionally refers to situations where (fine-tuned) systems with many degenerate ground states have their degeneracy lifted by fluctuations. Conventional “order by disorder” decreases the number of degenerate states, and thus in some sense actually reduces the amount of order. By contrast, the dynamics considered in this work uses fluctuations to increase the number of steady states, thereby increasing the amount of order. , and future work [18] will show how fluctuations beyond MF stabilize the ordered phase.
@@
To further explore this phenomenon, consider updates where each site has a probability $\alpha\in[0,1]$ of being updated at each time step. At small enough $p$, each of the studied memories undergoes a percolation transition at an intermediate update fraction $\alpha_{c}$ between a robust memory phase (at small $\alpha$), and a trivial phase dominated by $1$s (at large $\alpha$); this phenomenon is somewhat similar to the synchronicity transitions that occur in other (non-robust) CA, such as Conway’s game of life [27, 28]. As shown in Fig. 2, the spatial spin patterns produced across this transition can be quite rich.
Figure 3: Top: Probability that a rule remains a memory $P_{\sf mem}$ as a function of the number of “mutations” made to its rule table; each data point is based on sampling 5000 randomly mutated rules. Bottom: Hamming distances between pairs of distinct rule tables for the learned NCA and majority-vote memories. Each point is the distance between one rule from category A and one from category B.
@@
The phenomena discovered in this work indicate that the physics of non-equilibrium memories is quite diverse, and far from fully understood. Immediate future research directions include generalizing the NCA framework to $\mathbb{Z}_{N}$ and $U(1)$ memories, investigating the phase transitions that occur when the memory is lost, and theoretically characterizing the distinct universality classes of local error-correcting dynamics.
While our focus in this work has been entirely on classical memories, machine learning methods can also be used to discover local error-correcting dynamics for quantum memories. The simplest examples are closely related to one-dimensional classical memories, which remain poorly understood, although in this context, the (reasonable) assumption of noiseless classical error-correcting hardware makes the task easier. Future work will describe this in detail [29].
We close with some rather speculative remarks. The systems studied in this work are capable of “homeostasis”: the robust protection of (here, a single bit of) information against degradation by noise. It is natural to wonder about mechanisms by which richer behavior, like nontrivial information processing and environmental adaptation, can robustly emerge in locally-interacting systems. It is well known that simple CA rules can display remarkably life-like emergent behavior [30, 31, 8], but in none of these cases has the problem of noise resilience been carefully addressed, and an outstanding problem for theory is to identify the complexity threshold that must be crossed in order for CA to exhibit complex and robust emergent behavior. The constructions of Gacs [2, 32]—whose design calls to mind the genetic code of an elaborate artificial organism—provide extremely complicated starting points, but there is clearly much more to be done. We expect the NCA approach developed in the present work to be useful in this regard.
@@ Acknowledgments:
We thank Shankar Balasubramanian, Sarang Gopalakrishnan, Hsin-Yuan Huang, Vedika Khemani, Yaodong Li, Tibor Rakovszky, and Mike Zaletel for discussions and feedback.
A.B. is supported by the Kortschak Scholars Program. E.L. is supported by a Miller Research Fellowship. N.S. is supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
Author contributions:
A.B., E.L., and N.S. conceived the idea of using NCA to discover robust memories. A.B., E.P., and N.S. designed and implemented the numerics, E.P. built the interactive demo, and E.L. performed the theoretical calculations.